# Intuitionism with Truth Tables: A Decision Procedure for IPL Based on RNmatrices

### v2

@@ 4.2.3 Soundness and completeness of the intuitionistic truth tables
Suppose first that $\Lambda=For(\Omega)$. By induction on $n\in\omega$, it will be shown that $\tilde{v}_{p}\in\mathcal{L}^{IPL}_{n}$ for every $n$. By Definition 4.25 of partial valuation in $\mathcal{M}_{IPL}$, $\tilde{v}_{p}\in Val_{+}(\mathcal{M}_{IPL})={\mathcal{L}}_{0}^{IPL}$. Suppose that $\tilde{v}_{p}\in\mathcal{L}^{IPL}_{n}$ for a given $n\geq 0$, and let $\alpha\in For(\Omega)_{n+1}$ such that $\tilde{v}_{p}(\alpha)=\mathbf{U}$. By Definition 4.30 of intuitionistic partial’ level valuation, there exists a level valuation $w$ in $\mathcal{L}_{IPL}$ (and so, $w\in\mathcal{L}^{IPL}_{n}$) such that $\mathsf{P}_{\Lambda}^{i}(\tilde{v}_{p},w,\alpha)$. Since $\Lambda=For(\Omega)$, this means that $\tilde{v}_{p}\in\mathcal{L}^{IPL}_{n+1}$. From this, $\tilde{v}_{p}\in\mathcal{L}_{IPL}$ and the result clearly holds.
@@
Suppose now that $\Lambda\neq\emptyset$ is a proper subset of $For(\Omega)$. Consider an enumeration of $For(\Omega)$ as in Remark 4.33. The valuation $v$ will be defined by induction on the complexity $n$ of the formulas. Observe that $\Lambda\cup\mathcal{P}\in CS(\Omega)$ and, for every $n\in\omega$, $\bar{\Lambda}_{n}\in CS(\Omega)$ for $\bar{\Lambda}_{n}:=\Lambda\cup For(\Omega)_{n}$. In addition, for every $m\in\omega^{2}$ and $n\geq 0$, if $\alpha_{m}$ is such that $\textsf{c}(\alpha_{m})=n+1$, then $\bar{\Lambda}_{n}^{m}\in CS(\Sigma)$, where $\bar{\Lambda}_{n}^{m}=\bar{\Lambda}_{n}\cup\{\alpha_{m}\}$. Clearly $\bar{\Lambda}_{n+1}=\cup_{k\in\omega}\bar{\Lambda}_{n}^{\omega\cdot(n+1)+k}$.
@@
Base $n=0$: Define $v(\alpha)=\tilde{v}_{p}(\alpha)$ if $\alpha\in\Lambda$, and $v(p)=\mathbf{F}$ if $p$ is a propositional variable which does not belong to $\Lambda$. Clearly, $v\in PLV^{\prime}(\Lambda\cup\mathcal{P})$ up to this stage.
@@
Inductive step: Assume that $v(\alpha)$ was defined for every $\alpha\in\bar{\Lambda}_{n}$ such that $v(\alpha)=\tilde{v}_{p}(\alpha)$ if $\alpha\in\Lambda$, and $v\in iPLV^{\prime}(\bar{\Lambda}_{n})$ up to this stage, for a given $n\geq 0$ (Induction Hypothesis, IH). Let $\alpha=\alpha_{m}$ (for $m\in I_{n+1}$) such that $\alpha\not\in\Lambda$ and $\textsf{c}(\alpha)=n+1$. Observe that $v(\beta)$ was already defined, for every strict subformula $\beta$ of $\alpha$. Since $v\in iPLV^{\prime}(\bar{\Lambda}_{n})$ then, for every $\delta\in\bar{\Lambda}_{n}$ such that $v(\delta)=\mathbf{U}$ there exists $w\in\mathcal{L}_{IPL}$ such that $\mathsf{P}_{\bar{\Lambda}_{n}}^{i}(v,w,\delta)$. Such a $w$ will be denoted by $w^{\delta}$ (observe that it is possible to have more that one $w^{\delta}$ for each $\delta$, and it is possible to have $w^{\delta}=w^{\gamma}$ for $\delta\neq\gamma$). Recall that $\bar{\Lambda}_{n}^{m}=\bar{\Lambda}_{n}\cup\{\alpha_{m}\}$ is closed under subformulas. There are three cases to analyze:
@@
(1) $\alpha=\neg\beta$ such that $\neg^{\mathcal{A}_{IPL}}v(\beta)=\{\mathbf{U},\mathbf{T}\}$, or $\alpha=\beta\to\gamma$ such that $\to^{\mathcal{A}_{IPL}}(v(\beta),v(\gamma))=\{\mathbf{U},\mathbf{T}\}$. There are three subcases to analyze:
@@
(1.1) There is $\delta\in\bar{\Lambda}_{n}$ with $v(\delta)=\mathbf{U}$ and there exists some $w^{\delta}$ such that $w^{\delta}(\alpha)=\mathbf{F}$. Define $v(\alpha)=\mathbf{U}$. Hence $v\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m})$ such that $w^{\alpha}=w^{\delta}$.
@@
(1.2) There is $\delta\in\bar{\Lambda}_{n}$ with $v(\delta)=\mathbf{U}$ and there exists some $w^{\delta}$ such that $w^{\delta}(\alpha)=\mathbf{U}$. By Lemma 4.32, $v^{\prime}=w^{\delta}_{|\bar{\Lambda}_{n}^{m}}\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m})$, hence there exists $w^{\prime\prime}\in\mathcal{L}_{IPL}$ such that $\mathsf{P}_{\bar{\Lambda}_{n}^{m}}^{i}(v^{\prime},w^{\prime\prime},\alpha)$. Define $v(\alpha)=\mathbf{U}$. Hence, $v\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m})$ such that $w^{\alpha}=w^{\prime\prime}$.
@@
(1.3) For every $\delta\in\bar{\Lambda}_{n}$ such that $v(\delta)=\mathbf{U}$, $w^{\delta}(\alpha)=\mathbf{T}$ for every $w^{\delta}$. Define $v(\alpha)=\mathbf{T}$. Clearly, $v\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m})$.
[…]
We have shown in cases (1)-(3) how to extend the domain of $v$ to the additional formula $\alpha_{m}$, for every $\alpha_{m}$ such that $\textsf{c}(\alpha_{m})=n+1$, showing that $v\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m})$. Consider now the simultaneous extension of $v$ with the values assigned to every $\alpha_{m}$ as prescribed above. We obtain so an extension of $v$ to $\bar{\Lambda}_{n+1}=\cup_{m}\bar{\Lambda}_{n}^{m}$. By the very definition of intuitionistic partial’ level valuations it follows that $v\in iPLV^{\prime}(\bar{\Lambda}_{n+1})$, and $v(\alpha)=\tilde{v}_{p}(\alpha)$ if $\alpha\in\Lambda$. This completes the inductive step. ∎
@@
Let $\tilde{v}_{p}\in iPLV^{\prime}(\Lambda)$. Then, $\tilde{v}_{p}\in iPV(\Lambda)$. Suppose that $\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=\mathbf{U}$. Then, there exists $w\in\mathcal{L}_{IPL}$ such that $\mathsf{P}_{\Lambda}^{i}(\tilde{v}_{p},w,\alpha)$. Let $\tilde{w}_{p}=w_{|\Lambda}$. By Lemma 4.32, $\tilde{w}_{p}\in iPLV^{\prime}(\Lambda)$ such that $\mathsf{P}_{\Lambda}^{i}(\tilde{v}_{p},\tilde{w}_{p},\alpha)$. Now, suppose that there exists $\tilde{w}^{\prime}_{p}\in iPLV^{\prime}(\Lambda)$ such that $\mathsf{P}_{\Lambda}^{i}(\tilde{v}_{p},\tilde{w}^{\prime}_{p},\alpha)$. By Lemma 4.34, there exists $v^{\prime}\in\mathcal{L}_{IPL}$ such that the restriction of $v^{\prime}$ to $\Lambda$ coincides with $\tilde{w}^{\prime}_{p}$. Hence, $\mathsf{P}_{\Lambda}^{i}(\tilde{v}_{p},v^{\prime},\alpha)$. This shows that
$\begin{array}[]{lll}iPLV^{\prime}(\Lambda)&=&\{\tilde{v}_{p}\in PV(\Lambda)\mid\forall\alpha\in\Lambda\big{(}\tilde{v}_{p}(\alpha)=\mathbf{U}\mbox{ implies that }\mathsf{P}_{\Lambda}^{i}(\tilde{v}_{p},\tilde{w}_{p},\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $\tilde{w}_{p}\in iPLV^{\prime}(\Lambda)$}\big{)}\}.\end{array}$ From this, being $iPV(\Lambda)$ finite, it follows that $iPLV(\Lambda)=iPLV^{\prime}(\Lambda)$. ∎
@@
Proposition 4.38.
Let $\Gamma\cup\{\alpha,\beta\}\subseteq For(\Omega)$. If $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha\to\beta$ then $\Gamma,\alpha\models_{\mathcal{R}(\mathcal{M}_{IPL})}\beta$.
Proof.
Assume that $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha\to\beta$, and let $v\in\mathcal{L}_{IPL}$ such that $v(\gamma)=v(\alpha)=\mathbf{T}$ for every $\gamma\in\Gamma$. By hypothesis, $v(\alpha\to\beta)=\mathbf{T}$. By definition of $\to^{\mathcal{A}_{IPL}}$, and given that $v\in Val(\mathcal{M}_{IPL})$, it follows from this that $v(\beta)=\mathbf{T}$. This shows that $\Gamma,\alpha\models_{\mathcal{R}(\mathcal{M}_{IPL})}\beta$. ∎
Theorem 4.39 (Soundness of $\mathcal{H}_{IPL}$ w.r.t. $\mathcal{R}(\mathcal{M}_{IPL})$).
For every $\Gamma$ and $\varphi$: $\Gamma\vdash_{IPL}\varphi$ implies that $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$.
Proof.
Since the consequence relation $\mathcal{H}_{IPL}$ of the Hilbert calculus for IPL is finitary and satisfies the deduction metatheorem, and taking into account Proposition 4.38, it is enough to prove the result for $\Gamma=\emptyset$. We prove first the following two facts:
Fact 1: If $\varphi$ is an instance of an axiom in $\mathcal{H}_{IPL}$ then $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$. Indeed, let $\varphi$ be an instance of an axiom in $\mathcal{H}_{IPL}$, and let $\Lambda$ be the set of subformulas of $\varphi$. Then, $\Lambda\in FCS(\Omega)$. It is routine to check that the truth-table for $\varphi$ over $\Lambda$ validates $\varphi$ (see Appendix A). That is: for every intuitionistic partial level valuation $\tilde{v}_{p}\in iPLV(\Lambda)$, $\tilde{v}_{p}(\varphi)=\mathbf{T}$. By Theorem 4.37 it follows that $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$.
Fact 2: If $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha$ and $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha\to\beta$ then $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\beta$. Indeed, assume that $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha$ and $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha\to\beta$. If $v\in\mathcal{L}_{IPL}$ then $v(\alpha)=v(\alpha\to\beta)=\mathbf{T}$. By definition of $\to^{\mathcal{A}_{IPL}}$, and by the fact that $v\in Val(\mathcal{M}_{IPL})$, it follows that $v(\beta)=\mathbf{T}$. This means that $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\beta$.
Now, suppose that $\vdash_{IPL}\varphi$. By induction on the length of a derivation of $\varphi$ in $\mathcal{H}_{IPL}$, and by Fact 1 and Fact 2 it is immediate to prove that $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$. ∎
Corollary 4.39.1 (Finitariness for $\mathcal{R}(\mathcal{M}_{IPL})$).
Proof.
It follows from Theorems 4.23, 4.39 and the fact that $\mathcal{H}_{IPL}$ is finitary. ∎
Let $\Gamma\cup\{\alpha,\beta\}\subseteq For(\Omega)$. If $\Gamma,\alpha\models_{\mathcal{R}(\mathcal{M}_{IPL})}\beta$ then $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha\to\beta$.
Proof.
It follows from Theorems 4.23, 4.39 and the fact that $\mathcal{H}_{IPL}$ satisfies the deduction metatheorem. ∎
Corollary 4.39.3 (Deduction-detachment metatheorem for intuitionistic truth-tables).
@@
It follows from Theorem 4.37, Proposition 4.38 and Corollary 4.39.2. ∎
It is worth noting that the latter result is useful to reduce the complexity of the truth-tables: instead of generating the (reduced) truth table for a formula $\delta=\beta_{1}\to(\beta_{2}\to(\ldots\to(\beta_{n}\to\varphi)\ldots))$ to be tested as a tautology, it suffices generating the (reduced) truth-table for $\{\beta_{1},\ldots,\beta_{n},\varphi\}$ and then determine if, in every row such that $\beta_{1},\ldots,\beta_{n}$ receive simultaneously the value $\mathbf{T}$, $\varphi$ also receives the value $\mathbf{T}$. This reduces the number of rows required to test $\delta$.
Corollary 4.39.4 (Soundness and completeness of $\mathcal{H}_{IPL}$ w.r.t. intuitionistic truth-tables).
@@
It follows from Theorems 4.39, 4.23 and 4.37. ∎
The table is never empty.
Lemma 4.40.
Let $\Delta\in CS(\Sigma)$ and $\beta\in For(\Sigma)$. If $v_{p}:\Delta\to\mathcal{V}$ is $iPLV$ and $sub(\beta)\subseteq\Delta$, then $v_{p}$ has at least one extension $v^{\prime}_{p}:\Delta\cup\{\beta\}\to\mathcal{V}$ such that $v^{\prime}_{p}$ is $iPLV$.
Proof.
By hypothesis, $v_{p}$ is iPLV. Then, by analiticity of $\mathcal{R}(\mathcal{M}_{IPL})$, there is a level valuation $v:For(\Sigma)\to\mathcal{V}$ such that $v_{p}=v_{\mid\Delta}$, i.e, for every $\alpha\in\Delta$, $v(\alpha)=v_{p}(\alpha)$. More then that, $v$ is defined for $\beta$. Now, given that $sub(\beta)\subseteq\Delta$ and $\Delta\in CS(\Sigma)$, then $\Delta\cup\{\beta\}\in CS(\Sigma)$ and, by co-analiticity of $\mathcal{R}(\mathcal{M}_{IPL})$, $v^{\prime}_{p}=v_{\mid\Delta\cup\{\beta\}}$ is iPLV. ∎
5 Complexity
Given a non-deterministic matrix $\mathcal{M}=\langle\mathcal{V},\mathcal{D},\mathcal{O}\rangle$ defined over a language $\mathcal{L}$ and a formula $\alpha\in\mathcal{L}$, we can estimate the total number of rows of the final truth table for $\alpha$ via the following upper bound.
Let $sub(\alpha)\subseteq\mathcal{L}$ be the set of all sub-formulas of $\alpha$, $atoms(\alpha)\subseteq sub(\alpha)$ the set of sub-formulas whose complexity is 00, and $\kappa\geq 1$ the branching factor. Then, the maximum number of rows of the truth table for $\alpha$ in $\mathcal{M}$ is given by
Definition 5.1.
$ub_{\mathcal{M}}(\alpha)=\kappa^{(|sub(\alpha)|-|atoms(\alpha)|)}\times|\mathcal{V}|^{|atoms(\alpha)|}$ (1)
In 5.1, we assume that the branching factor is the maximum degree of non-determinism. More formally,
Definition 5.2.

### v4

@@ 9.2 Soundness and completeness of the intuitionistic truth-tables
Suppose first that $\Lambda=For(\Omega)$. By induction on $n\in\omega$, it will be shown that $\tilde{v}_{p}\in\mathcal{L}^{IPL}_{n}$ for every $n$. By Definition 9.2 of partial valuation in $\mathcal{M}_{IPL}$, $\tilde{v}_{p}\in Val^{nd}_{+}(\mathcal{M}_{IPL})={\mathcal{L}}_{0}^{IPL}$. Suppose that $\tilde{v}_{p}\in\mathcal{L}^{IPL}_{n}$ for a given $n\geq 0$, and let $\alpha\in For(\Omega)$ such that $\tilde{v}_{p}(\alpha)=\mathbf{U}$. By Definition 9.11 of intuitionistic partial’ level valuation, there exists a level valuation $w$ in $\mathcal{L}_{IPL}$ (and so, $w\in\mathcal{L}^{IPL}_{n}$) such that $\mathsf{P}_{\Lambda}(\tilde{v}_{p},w,\alpha)$. Since $\Lambda=For(\Omega)$, this means that $\tilde{v}_{p}\in\mathcal{L}^{IPL}_{n+1}$. From this, $\tilde{v}_{p}\in\mathcal{L}_{IPL}$ and the result clearly holds by taking $v:=\tilde{v}_{p}$.
@@
Suppose now that $\Lambda\neq\emptyset$ is a proper subset of $For(\Omega)$. Consider an enumeration of $For(\Omega)$ as in Remark 9.15. The valuation $v$ will be defined by induction on the complexity $n$ of the formulas. More precisely, with terminology as in Definition 9.16 and from the observations in Remark 9.17, an extension $v_{n}^{m}$ of $\tilde{v}_{p}$ (i.e., $\tilde{v}_{p}\subseteq v_{n}^{m}$) will be defined for every $(n,w)\in\omega\times\omega$ such that $v_{n}^{m}\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m})$ and $v_{i}^{j}\subseteq v_{n}^{m}$ if $(i,j)\leq(n,m)$, where: $(i,j)\leq(n,m)$ iff $i\leq n$ or $i=n$ and $j\leq m$ (and where $\omega\times\omega$ denotes, as usual, the Cartesian product of $\omega$ with itself).
@@
Base $n=m=0$: Observe that $\bar{\Lambda}_{0}=\bar{\Lambda}_{0}^{0}=\Lambda\cup\mathcal{P}$. Define $v_{0}^{0}(\alpha)=\tilde{v}_{p}(\alpha)$ if $\alpha\in\Lambda$, and $v_{0}^{0}(p)=\mathbf{F}$ if $p$ is a propositional variable which does not belong to $\Lambda$. Clearly, $\tilde{v}_{p}\subseteq v_{0}^{0}$ and $v_{0}^{0}\in iPLV^{\prime}(\bar{\Lambda}_{0})=iPLV^{\prime}(\bar{\Lambda}_{0}^{0})$. Let $v_{0}:=v_{0}^{0}$.
Inductive step: Assume that $v_{i}^{j}(\alpha)$ was defined for every $\alpha\in\bar{\Lambda}_{i}^{j}$ such that $v_{i}^{j}(\alpha)=\tilde{v}_{p}(\alpha)$ if $\alpha\in\Lambda$, $v_{i}^{j}\in iPLV^{\prime}(\bar{\Lambda}_{i}^{j})$, $v_{i}^{j}\subseteq v_{k}^{r}$ if $(i,j)\leq(k,r)\leq(n,m)$, for given $n,m\geq 0$ (Induction Hypothesis, IH). Now it will be shown how to extend $v_{n}^{m}$ to a function $v_{n}^{m+1}$ with domain $\bar{\Lambda}_{n}^{m+1}=\bar{\Lambda}_{n}^{m}\cup\{\alpha_{\omega\cdot(n+1)+m}\}$. That is, it will be shown how to define a value for $\alpha_{\omega\cdot(n+1)+m}$ while preserving the values assigned by $v_{n}^{m}$, in such a manner that $v_{n}^{m+1}\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m+1})$. To start with, define $v_{n}^{m+1}(\beta)=v_{n}^{m}(\beta)$ for every $\beta\in\bar{\Lambda}_{n}^{m}$.
@@
Let $\alpha=\alpha_{\omega\cdot(n+1)+m}$. If $\alpha\in\Lambda$ then $\bar{\Lambda}_{n}^{m+1}=\bar{\Lambda}_{n}^{m}$. Then $v_{n}^{m+1}:=v_{n}^{m}$, and the required hypothesis are obviously satisfied by $v_{n}^{m+1}$. Suppose now that $\alpha\not\in\Lambda$. Observe that $v_{n}^{m+1}(\beta)=v_{n}^{m}(\beta)$ was already defined, for every strict subformula $\beta$ of $\alpha$. Since $v_{n}^{m}\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m})$ then, for every $\delta\in\bar{\Lambda}_{n}^{m}$ such that $v_{n}^{m}(\delta)=\mathbf{U}$ there exists $w\in\mathcal{L}_{IPL}$ such that $\mathsf{P}_{\bar{\Lambda}_{n}^{m}}(v_{n}^{m},w,\delta)$. Such a $w$ will be denoted by $w^{\delta}$ (observe that it is possible to have more than one $w^{\delta}$ for each $\delta$, and it is possible to have $w^{\delta}=w^{\gamma}$ for $\delta\neq\gamma$). Recall that $\bar{\Lambda}_{n}^{m}$ is closed under subformulas. There are three cases to analyze:
@@
(1) $\alpha=\neg\beta$ such that $\neg^{IPL}(v_{n}^{m}(\beta))=\{\mathbf{U},\mathbf{T}\}$, or $\alpha=\beta\to\gamma$ such that $\to^{IPL}(v_{n}^{m}(\beta),v_{n}^{m}(\gamma))=\{\mathbf{U},\mathbf{T}\}$. There are three subcases to analyze:
@@
(1.1) There is $\delta\in\bar{\Lambda}_{n}^{m}$ with $v_{n}^{m}(\delta)=\mathbf{U}$ and there exists some $w^{\delta}$ such that $w^{\delta}(\alpha)=\mathbf{F}$. Define $v_{n}^{m+1}(\alpha)=\mathbf{U}$. Hence $v_{n}^{m+1}\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m+1})$ such that $w^{\alpha}=w^{\delta}$.
@@
(1.2) There is $\delta\in\bar{\Lambda}_{n}^{m}$ with $v_{n}^{m}(\delta)=\mathbf{U}$ and there exists some $w^{\delta}$ such that $w^{\delta}(\alpha)=\mathbf{U}$. By Proposition 8.21, there exists $w^{\prime\prime}\in\mathcal{L}_{IPL}$ such that $\mathsf{P}_{For(\Omega)}(w^{\delta},w^{\prime\prime},\alpha)$. Define $v_{n}^{m+1}(\alpha)=\mathbf{U}$. Hence, $v_{n}^{m+1}\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m+1})$ such that $w^{\alpha}=w^{\prime\prime}$.
@@
(1.3) For every $\delta\in\bar{\Lambda}_{n}^{m}$ such that $v_{n}^{m}(\delta)=\mathbf{U}$, $w^{\delta}(\alpha)=\mathbf{T}$ for every $w^{\delta}$. Define $v_{n}^{m+1}(\alpha)=\mathbf{T}$. Clearly, $v_{n}^{m+1}\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m+1})$.
(2) $\alpha=\neg\beta$ such that $\neg^{IPL}(v_{n}^{m}(\beta))=\{\mathbf{F}\}$, or $\alpha=\beta\#\gamma$ such that $^{IPL}(v_{n}^{m}(\beta),v_{n}^{m}(\gamma))=\{\mathbf{F}\}$ (for some $#\in\{\vee,\land,\to\}$). Define $v_{n}^{m+1}(\alpha)=\mathbf{F}$. It is clear that $v_{n}^{m+1}\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m+1})$.
(3) $\alpha=\beta\#\gamma$ such that $^{IPL}(v_{n}^{m}(\beta),v_{n}^{m}(\gamma))=\{\mathbf{T}\}$ (for some $#\in\{\vee,\land,\to\}$). Define $v_{n}^{m+1}(\alpha)=\mathbf{T}$. It is clear that $v_{n}^{m+1}\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m+1})$.
We have shown in cases (1)-(3) how to extend the domain of $v_{n}^{m}$ to the additional formula $\alpha_{\omega\cdot(n+1)+m}\in\bar{\Lambda}_{n}^{m+1}$, showing that the resulting function $v_{n}^{m+1}$ is in $iPLV^{\prime}(\bar{\Lambda}_{n}^{m+1})$, and $\tilde{v}_{p}\subseteq v_{i}^{j}\subseteq v_{k}^{r}$, for every $(i,j)\leq(k,r)\leq(n,m+1)$. That is, we show how to increase the superscript $m$. In order to increase the subscript $n$, let $v_{n+1}=v_{n+1}^{0}:=\bigcup_{m\in\omega}v_{n}^{m}$. By the procedure described above, it is immediate to see that $v_{n+1}^{0}$ is in $iPLV^{\prime}(\bigcup_{m\in\omega}\bar{\Lambda}_{n}^{m})=iPLV^{\prime}(\bar{\Lambda}_{n+1})=iPLV^{\prime}(\bar{\Lambda}_{n+1}^{0})$, and $\tilde{v}_{p}\subseteq v_{i}^{j}\subseteq v_{k}^{r}$, for every $(i,j)\leq(k,r)\leq(n+1,0)$. By repeating the process, it is possible to define $v_{n}^{m}$ for every $(n,m)\in\omega\times\omega$ with the required properties. Finally, let $v:=\bigcup_{n\in\omega}v_{n}$. From the manner in which the construction was made, it follows that $v$ is in $iPLV^{\prime}(\bigcup_{n\in\omega}\bar{\Lambda}_{n})=iPLV^{\prime}(For(\Omega))$, and $\tilde{v}_{p}\subseteq v$. By the first part of the proof, we conclude that $v\in\mathcal{L}_{IPL}$. ∎
@@
Let $\tilde{v}_{p}\in iPLV^{\prime}(\Lambda)$. Then, $\tilde{v}_{p}\in iPV(\Lambda)$. Suppose that $\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=\mathbf{U}$. Then, there exists $w\in\mathcal{L}_{IPL}$ such that $\mathsf{P}_{\Lambda}(\tilde{v}_{p},w,\alpha)$. Let $\tilde{w}_{p}=w_{|\Lambda}$. By Lemma 9.13, $\tilde{w}_{p}\in iPLV^{\prime}(\Lambda)$ such that $\mathsf{P}_{\Lambda}(\tilde{v}_{p},\tilde{w}_{p},\alpha)$. Now, suppose that there exists $\tilde{w}^{\prime}_{p}\in iPLV^{\prime}(\Lambda)$ such that $\mathsf{P}_{\Lambda}(\tilde{v}_{p},\tilde{w}^{\prime}_{p},\alpha)$. By Lemma 9.18, there exists $v^{\prime}\in\mathcal{L}_{IPL}$ such that the restriction of $v^{\prime}$ to $\Lambda$ coincides with $\tilde{w}^{\prime}_{p}$. Hence, $\mathsf{P}_{\Lambda}(\tilde{v}_{p},v^{\prime},\alpha)$. This shows that
$\begin{array}[]{lll}iPLV^{\prime}(\Lambda)&=&\{\tilde{v}_{p}\in iPV(\Lambda)\mid\forall\alpha\in\Lambda\big(\tilde{v}_{p}(\alpha)=\mathbf{U}\mbox{ implies that }\mathsf{P}_{\Lambda}(\tilde{v}_{p},\tilde{w}_{p},\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $\tilde{w}_{p}\in iPLV^{\prime}(\Lambda)$}\big)\}.\end{array}$ From this, it follows that $iPLV(\Lambda)=iPLV^{\prime}(\Lambda)$, by Proposition 9.7. ∎
Remark 9.20.
In Proposition 9.7 the set $iPLV(\Lambda)$ was characterized constructively, as the output of an algorithm (Algorithm 1). In this subsection, we define an alternative subset $iPLV^{\prime}(\Lambda)\subseteq iPV(\Lambda)$ in a predicative way. This subset connects the RNmatrix semantics for $\mathbf{IPL}$ (based on level valuations) to the truth-table procedure given by $iPLV(\Lambda)$. Analyticity and co-analyticity of $iPLV^{\prime}(\Lambda)$ for level valuations allow us to show that, for a finite $\Lambda$, $iPLV^{\prime}(\Lambda)$ satisfies the condition of Remark 9.4. By uniqueness (Proposition 9.7), $iPLV^{\prime}(\Lambda)$ must coincide with $iPLV(\Lambda)$, i.e., with the truth-table generated by $\Lambda$. But then we obtain analyticity and co-analyticity of $iPLV(\Lambda)$ for level valuations, which implies the soundness and completeness of truth-tables for $\mathbf{IPL}$, see Theorem 9.22.
@@
Finally, we establish the correspondence between the restricted Nmatrix for $\mathbf{IPL}$ and the intuitionistic truth-tables:
@@
Corollary 9.22.1 (Deduction metatheorem for intuitionistic truth-tables).
@@
It follows from Theorem 9.22 and Corollary 8.19.2. ∎
It is worth noting that the latter result is useful to reduce the complexity of the truth-tables: instead of generating the (reduced) truth-table for a formula $\delta=\beta_{1}\to(\beta_{2}\to(\ldots\to(\beta_{n}\to\varphi)\ldots))$ to be tested as a tautology, it suffices generating the (reduced) truth-table for $\{\beta_{1},\ldots,\beta_{n},\varphi\}$ and then determine if, in every row such that $\beta_{1},\ldots,\beta_{n}$ receive simultaneously the value $\mathbf{T}$, $\varphi$ also receives the value $\mathbf{T}$. This reduces the number of rows required to test $\delta$.
Finally, we arrive to the main result of the paper, recalling that LI is a standard sequent calculus for $\mathbf{IPL}$:
Theorem 9.23 (Soundness and completeness of LI w.r.t. intuitionistic truth-tables).
@@
It follows from Theorems 8.12, 8.18 and 9.22. ∎
The latter result shows that the procedure of constructing (non-deterministic) truth-tables in Definition 9.9 determines a decision method for propositional intuitionistic logic $\mathbf{IPL}$.
10 Some considerations on complexity
Unlike deterministic truth-tables, the size of non-deterministic truth-tables cannot be determined solely by the number of atoms. This is because, in principle, each subformula may (or may not) create new rows. Thus, we must also consider the number and shape of every evaluated subformula.
A workaround for this problem is to establish size limits using boundaries. More formally, we can define a general upper bound for non-deterministic matrices by identifying their branching factor as follows.
Definition 10.1 (Branching factor).